# Linying Lv

### v1

@@ 3.2.5 Comparison with Analyst Report Sentiment
Huang et al. (2014) emphasize the importance of analyst report text by extracting tone using a Naive Bayes machine learning approach. My study extends this research by incorporating more advanced techniques, including BERT-based tone measures and text embeddings.
@@
I construct report-level tone measures using both the Naive Bayes approach and BERT classification. Following Huang et al. (2014), I manually label 10,000 sentences as a training sample for both models. Each sentence is categorized into positive, neutral, or negative using these pre-trained models, then average the tone at the report level to derive $Tone\_NB$ and $Tone\_BERT$ measures. I also construct tone measures separately for Income Statement topics and non-Income Statement topics.
@@
Table 6 presents a comparison of performance between text tone measures and text embeddings. Panel A summarizes the OLS regression of three-day $CAR$ on the tone measures. The coefficient estimate for $Tone\_NB$ in column (1) is 0.019, comparable to the 0.0208 reported by Huang et al. (2014). A one standard deviation increase in the favorableness of the text tone results in an additional $CAR$ of 49 basis points, which is comparable with the 41 basis points increase in two-day abnormal return in Huang et al. (2014). Column (2) summarizes the results of regressing $CAR$ on tone measures extracted within Income Statement topics and non-Income Statement topics. Both coefficients of tone are significantly positive. Column (3) reports regression results with firm-fixed and year-fixed effects, consistent with the findings in column (2).
@@
Columns (4)-(6) present regression results with BERT-based tone measures as explanatory variables. The coefficient estimates’ magnitude increases dramatically, along with significance levels. Economically, a one standard deviation increase in BERT-based text tone favorableness results in an additional three-day $CAR$ of 86 basis points. When both tone measures are included in an OLS regression in column (7), Naive Bayes-based tone measure coefficients become insignificant, suggesting BERT-based measures capture sentiment more accurately and effectively.
@@
Panel B of Table 6 compares the $R^{2}_{\mathrm{OOS}}$ of extracted tone measures and embedding representations. The $R^{2}_{\mathrm{OOS}}$ of tone inputs ranges from 0.05% to 3.68% in the first four rows. The $R^{2}_{\mathrm{OOS}}$ increases from 9.49% to 12.27% after incorporating analyst revision and text embeddings, confirming the superior performance of state-of-the-art LLMs across various NLP tasks. The bottom two rows of Panel B report comparisons of $R^{2}_{\mathrm{OOS}}$ and DM tests between models with “Tone + Rev + text” versus “Rev + text.” The $R^{2}_{\mathrm{OOS}}$ difference between these models is not statistically distinguishable from zero for both Naive Bayes-based and BERT-based tone measures.
@@
In general, the empirical results in Table 6 support two key findings regarding the superiority of LLMs compared to conventional NLPs. First, overall report text tone measures extracted using small-size LLMs are more explanatory for contemporaneous stock returns than those extracted using the Naive Bayes algorithm. Second, tone measures (extracted with both Naive Bayes and BERT) add no incremental value to text embeddings, suggesting that LLMs provide more comprehensive contextualized representations.
@@ 3.2.6 Information Content with Alternative Machine Learning Models
In this Section, I examine the information content using three alternative machine learning algorithms, applying the same processes as in Ridge regressions. Each model uses full-context report embeddings as inputs, with the target variable being the three-day $CAR$. To ensure out-of-sample evaluation, I train the models yearly and estimate for the following year’s sample. The $R^{2}_{\mathrm{OOS}}$ is calculated by averaging the estimated and realized $CAR$ values. Additionally, I perform pairwise Diebold-Mariano tests using a zero benchmark to assess predictive accuracy.
I begin with PLS regression, summarized in Table 7. PLS is notable for extracting the most relevant common components from sentiment proxies that align with the target variable (Huang et al., 2015). If there is a strong common component in the text embeddings that is informative of stock returns, the PLS regression architecture should effectively capture this relevant information. Examining the model performance, I observe that PLS regressions produce an overall $R^{2}_{\mathrm{OOS}}$ of 8.26%. This measure of information content ranges from 2.14% in the 2021 sample to 12.23% in the 2015 sample, showing a similar pattern across years as seen in Ridge models. Diebold-Mariano tests indicate the $R^{2}_{\mathrm{OOS}}$ is significant at the 1% level across all periods, demonstrating strong evidence of information content in the analyst report text.
@@
Next, I replicate the analyses with the XGBoost algorithm. Gradient boosting combines weak estimations to form a strong estimation by constructing sequential trees. Each tree is responsible for training a small sample to fit the residuals of the previous trees and updating the model by learning new information. Therefore, if the high-dimensional feature spaces of text embeddings contain segmented information with discrete importance, XGBoost can potentially identify the most informative aspects of the analyst report embeddings for explaining contemporaneous stock returns. Compared to Ridge and PLS models, XGBoost is capable of capturing complex nonlinear relationships between analyst report embeddings and the target variable. As summarized in Table 7, the XGBoost algorithm underperforms with an overall $R^{2}_{\mathrm{OOS}}$ of 5.97% (DM t-stat = 6.76). The relatively poor performance could be attributed to the inherent complexity of language, where information is distributed across the embedding space rather than concentrated in specific dimensions.
@@
Neural Networks, forming the core of LLMs and the transformer architecture, have shown significant advancements in various NLP tasks (Cao et al., 2024; Li et al., 2023). In a Neural Network, nodes are interconnected, mimicking the neuron structure of the human brain. Since embeddings are encoded as nodes within the Neural Network layers, I expect this architecture to capture long-range relationships between words and the complex contextual information embedded in the text. I construct the Neural Network estimations as an ensemble of five models for each NN configuration. As shown in Table 7, the $R^{2}_{\mathrm{OOS}}$ of Neural Network models ranges from 10.42% to 12.12%, achieving the highest performance among all tested ML algorithms. Performance tends to improve with network depth, aligning with the virtue of complexity principle proposed by Kelly et al. (2022).
@@
The experiments with machine learning models yield noteworthy findings. First, the information content of report text, as indicated by $R^{2}_{\mathrm{OOS}}$, achieves around 10% in certain models, which is significant considering the low signal-to-noise ratio in stock returns. Second, the superior performance of deep learning architectures emphasizes the importance of capturing contextual information and word interdependencies for textual analyses.
@@ 3.3 Information Value of Analyst Reports
How can the statistical measure $R^{2}_{\mathrm{OOS}}$ be translated into economic terms? In this section, I quantify the information value of analyst reports using the measure introduced in Section 2.3. This measure reflects the profit derived from a reduction in price uncertainty due to an informative signal while considering price impact. In other words, both greater information content in analyst reports and lower price impact will lead to higher estimated information value.
@@
These patterns indicate that the information value of analyst reports is dynamic, and influenced by market conditions and external events. The relationship between combined and text-only information value fluctuates over time, highlighting the importance of considering both numerical and textual components when analyzing analyst reports.
In the Online Appendix, Table A7 provides the summary statistics of alternative estimates for information value. Using high-frequency measures of realized return volatility and different algorithms for signing trades as buys or sells, the estimates for information value from analyst reports range from \$0.42 million to \$0.64 million. Figure A8 illustrates the time trends of these three alternative estimates, revealing consistent patterns in the evolution of information value over time. These alternative estimates support my main findings, reinforcing the substantial value of analyst information and the robustness of the approach.
@@ 3.3.1 The Value of Analyst Information in the Cross-section of Stocks
It is likely that the information value of sell-side analysts will vary in the cross-section of stocks. Specifically, the costs of information processing and dissemination may differ by firm size (Zeghal, 1984). For example, analyst opinions on smaller stocks may trigger larger market reactions, as these stocks generally have less publicly available information (Stickel, 1995). I plot the scatter of information value and market capitalization by stock in Figure A9. There is a general upward trend, indicating that stocks with higher market equity tend to have higher information values. However, the relationship is not strictly linear, as there are notable exceptions. I formally test the relationship by regressing the value of information on stock size, firm fixed effects, and year fixed effects. I also control for the book-to-market ratio in the regressions.
Table 9 shows the results. Stocks with larger market equity have significantly higher information value from analysts. Economically, column (4) means a 1% increase in market equity is associated with a 0.864% increase in analyst information value. Columns (5)-(8) reveal that the information value of report text is more sensitive to size compared to the combined numerical and textual information value. I further investigate whether the positive size elasticity of information value is driven by higher explained return volatility or lower price impact for large stocks. This decomposition can be expressed logarithmically as:
@@
Panel B of Table 9 demonstrates that the main reason behind the positive size elasticity of information value is the higher liquidity of large stocks. Overall, analyst information does not appear to explain the higher contemporaneous return volatility of large stocks, yet it remains more valuable due to the higher liquidity of these stocks.
@@ 3.3.2 The Value of Analyst Information in the Cross-section of Analysts
I next examine how the information value of analysts differs with herding behavior. Prior research has stated that analysts making bold forecasts incorporate their private information more completely and provide more relevant information to investors than those making herding forecasts (Clement and Tse, 2005).
@@
Table 10 presents the results. In columns (1)-(4) of Panel A, the coefficient estimates on $Bold$ are significantly positive, suggesting that analysts’ information is more strategically valuable when they provide bold forecasts. On average, information from bold reports is 29% more valuable than that from herding reports. The magnitude remains largely unchanged when testing with information value derived from report text. However, the insignificant coefficient estimates for $Brokersize$ suggest that the value of analyst information is not necessarily tied to the size of the brokerage firm.
The subsequent research question is whether the association between boldness and information value is driven by higher information content or lower price impact. In Panel B, I regress the two decompositions of information value on the boldness indicator. As shown in columns (1)-(4), information from analysts who issue bold forecasts explains higher return volatility, supporting the argument of Clement and Tse (2005) that analysts provide bold forecasts when they possess relevant private information. In contrast, the coefficient estimates in columns (5)-(8), which examine the relationship between boldness and price impact, are statistically insignificant. This finding suggests that while bold forecasts contain more valuable information, they do not necessarily lead to a higher price impact compared to herding forecasts.
@@ 3.3.3 The Value of Analyst Information around Earnings Announcements
Previous evidence underlines the information content of analyst reports in the first week following earnings announcements. To expand on this, I evaluate the information value of analyst reports issued promptly following earnings announcements and beyond that.
@@
To formally validate the observed pattern, I conduct a regression analysis of log information value on an indicator variable $Week$ which takes the value of one if the report is released within one week following earnings announcements and its interaction with earnings announcement day trading volume ($TradingVolume$). This model is motivated by prior literature suggesting that analysts’ interpretative role becomes more prominent with the complexity of financial accounting information (Chen et al., 2010). The trading volume serves as a proxy for investor disagreement, following Banerjee (2011) and Beckmann et al. (2024). The interaction term ($Week\times TradingVolume$) captures how post-earnings timing and investor disagreement jointly affect information value. The model includes stock and year-fixed effects to control for time-invariant stock characteristics and market-wide yearly trends.
@@
Panel B reveals significant increases in both log return variance and log price impact following earnings announcements, as evidenced by positive coefficients for the $Week$ variable. The interaction term ($Week\times TradingVolume$) is significantly positive in columns (1)-(4), indicating that explained return volatility rises with higher trading volumes during post-earnings periods. Notably, the increase in the information value of analyst reports following earnings announcements is primarily attributed to higher explained return volatility, rather than changes in price impact. These findings suggest that analyst reports provide more valuable insights in the immediate post-earnings announcement period, particularly when investor uncertainty is high. This increased value likely reflects analysts’ ability to interpret complex earnings information, helping investors make sense of the latest financial data during these critical periods.
@@ A. Additional Figures
Figure A6: Shapley Values for Topic Importance across Industries

### v2

@@ 3.2.5 Comparison with Analyst Report Sentiment
Huang et al. (2014) emphasize the information content of analyst report text by extracting tone using a Naive Bayes machine learning approach. I compare report-level tone measures derived from both Naive Bayes and BERT classification. Following Huang et al. (2014), I manually label 10,000 sentences as training data for both models. Each sentence is categorized as positive, neutral, or negative using these pre-trained models, and I average the tone at the report level to create $Tone_{N}B$ and $Tone_{B}ERT$ measures. I also construct tone measures separately for Income Statement topics and non-Income Statement topics.
@@
Table 6 presents a comparison of performance between text tone measures and text embeddings. Panel A summarizes the OLS regression of three-day $CAR$ on the tone measures. The coefficient estimate for $Tone\_NB$ in column (1) is 0.019, comparable to the 0.0208 reported by Huang et al. (2014). A one standard deviation increase in the favorableness of the text tone results in an additional $CAR$ of 49 basis points, comparable with the 41 basis points increase in two-day abnormal return reported in Huang et al. (2014). Column (2) summarizes the results of regressing $CAR$ on tone measures extracted within Income Statement topics and non-Income Statement topics. Both coefficients of tone are significantly positive. Regression results with firm-fixed and year-fixed effects are consistent, as reported in column (3).
@@
Columns (4)-(6) summarizes regression results with BERT-based tone measures as explanatory variables. The coefficient estimates’ magnitude increases dramatically, along with significance levels. Economically, a one standard deviation increase in BERT-based text tone favorableness results in an additional three-day $CAR$ of 86 basis points. When both tone measures are included in an OLS regression in column (7), coefficient estimates of Naive Bayes-based tone measure become insignificant, suggesting that BERT model captures sentiment more accurately and effectively.
@@
Panel B of Table 6 compares the $R^{2}_{\mathrm{OOS}}$ of tone measures and embeddings. The $R^{2}_{\mathrm{OOS}}$ of tone inputs ranges from 0.05% to 3.68% in the first four rows. The $R^{2}_{\mathrm{OOS}}$ improves to 9.49% and 12.27% after incorporating forecast revisions and text embeddings. The bottom two rows of Panel B report comparisons of $R^{2}_{\mathrm{OOS}}$ and DM tests between models with “Tone + Rev + text” versus “Rev + text.” The difference between the $R^{2}_{\mathrm{OOS}}$ of models with and without tone measures is not statistically distinguishable from zero.
@@
The empirical results in Table 6 highlight two key advantages of LLMs over conventional NLP methods. First, tone measures extracted using LLMs better explain contemporaneous stock returns than those from the Naive Bayes algorithm. Second, neither BERT nor Naive Bayes tone measures add incremental value beyond text embeddings, suggesting that LLMs capture richer contextual information than simple tone metrics.
@@ 3.2.6 Information Content with Alternative Machine Learning Models
In this Section, I examine the information content using three alternative machine learning algorithms, applying the same processes as in Ridge regressions. Each model uses full-context report embeddings as inputs, with the target variable being the three-day $CAR$.
Table 7 presents results from PLS regressions. PLS effectively extracts common components that correlate with the target variable. If text embeddings contain strong common components that predict stock returns, PLS regression should capture this information. The PLS regressions achieve an $R^{2}_{\mathrm{OOS}}$ of 8.26%, with year-to-year variations similar to those in Ridge models. DM tests confirm that the $R^{2}_{\mathrm{OOS}}$ is significant at the 1% level across all years.
@@
Next, I replicate the analyses with the XGBoost algorithm. Gradient boosting combines weak estimations to form a strong estimation by constructing sequential trees. Each tree is responsible for training a small sample to fit the residuals of the previous trees and updating the model by learning new information. Therefore, if the high-dimensional feature spaces of text embeddings contain segmented information with discrete importance, XGBoost can potentially identify the most informative aspects of the analyst report embeddings for explaining contemporaneous stock returns. Compared to Ridge and PLS models, XGBoost is capable of capturing complex nonlinear relationships between analyst report embeddings and the target variable. As summarized in Table 7, the XGBoost algorithm underperforms Ridge regressions with an overall $R^{2}_{\mathrm{OOS}}$ of 5.97% (DM t-stat = 6.76). The relatively poor performance could be attributed to the inherent complexity of language, where information is distributed across the embedding space rather than concentrated in specific dimensions.
@@
I then proceed with Neural Network models, which form the core of LLMs and the transformer architecture. In a Neural Network, nodes are interconnected, mimicking the neuron structure of the human brain. Since embeddings are encoded as nodes within the Neural Network layers, I expect this architecture to capture long-range relationships between words and the complex contextual information embedded in the text. I construct the Neural Network estimations as an ensemble of five models for each NN configuration. As shown in Table 7, the $R^{2}_{\mathrm{OOS}}$ of Neural Network models ranges from 10.42% to 12.12%, achieving the highest performance among all tested ML algorithms. Performance tends to improve with network depth, aligning with the virtue of complexity principle proposed by Kelly et al. (2022).
@@
The experiments with machine learning models yield noteworthy findings. First, the information content of report text, as indicated by $R^{2}_{\mathrm{OOS}}$, achieves around 10% in certain models, which is significant considering the low signal-to-noise ratio in stock returns. Second, the superior performance of deep learning architectures emphasizes the importance of contextual information and word interdependencies in textual analyses.
@@ 3.3 Information Value of Analyst Reports
How to translate statistical measure $R^{2}_{\mathrm{OOS}}$ into economic terms? In this section, I quantify the information value of analyst reports using the measure introduced in Section 2.3. This measure reflects the profit derived from a reduction in price uncertainty due to an informative signal, while accounting for price impact. In other words, both greater information content in analyst reports and lower price impact will lead to higher estimated information value.
@@
These patterns indicate that the information value of analyst reports is dynamic, and influenced by market conditions. The relationship between combined and text-only information value fluctuates over time, highlighting the importance of considering both numerical and textual components when analyzing analyst reports.
In the Online Appendix, Table A7 provides the summary statistics of alternative estimates for information value. Using high-frequency measures of realized return volatility and different algorithms for signing trades as buys or sells, the estimates for information value from analyst reports range from \$0.42 million to \$0.64 million. Figure A8 illustrates the time trends of these three alternative estimates, revealing consistent patterns in the evolution of information value over time.
@@ 3.3.1 The Value of Analyst Information in the Cross-section of Stocks
The information value of sell-side analysts likely vary in the cross-section of stocks. Specifically, the costs of information processing and dissemination may differ by firm size (Zeghal, 1984). For example, analyst opinions on smaller stocks may trigger larger market reactions, as these stocks generally have less publicly available information (Stickel, 1995). I plot the scatter of information value and market capitalization by stock in Figure A9. There is a general upward trend, indicating that stocks with higher market equity tend to have higher information values. However, the relationship is not strictly linear, as there are notable exceptions. I formally test the relationship by regressing the value of information on stock size, firm fixed effects, and year fixed effects. I also control for the book-to-market ratio in the regressions.
Table 9 shows the results. Stocks with larger market equity have significantly higher information value from analysts. Economically, results in column (4) indicate a 1% increase in market equity is associated with a 0.864% increase in analyst information value. Columns (5)-(8) reveal that the information value of report text is more sensitive to size compared to the combined numerical and textual information value. I further investigate whether the positive size elasticity of information value is driven by higher explained return volatility or lower price impact of large stocks. The decomposition can be expressed logarithmically as:
@@
Panel B of Table 9 shows that the higher information value for large stocks stems primarily from their greater liquidity. Analyst reports of large stocks do not explain higher contemporaneous return volatility, yet their value increases with stock size due to enhanced liquidity.
@@ 3.3.2 The Value of Analyst Information in the Cross-section of Analysts
Analysts who make bold forecasts incorporate their private information more completely and provide more relevant information to investors than those who herd in their forecasts (Clement and Tse, 2005). I examine how information value varies with analysts’ herding behavior.
@@
Table 10 presents the results. In columns (1)-(4) of Panel A, the coefficient estimates on $Bold$ are significantly positive, suggesting that analysts’ information is more strategically valuable when they provide bold forecasts. On average, information from bold reports is 29% more valuable than that from herding reports. This magnitude persists when examining the information value of report text alone. The insignificant coefficients for $Brokersize$ suggest that the value of analyst information is not related to brokerage firm size.
A subsequent research question is whether the relationship between analyst boldness and information value is driven by higher report informativeness or lower price impact. In Panel B, I regress the two decompositions of information value on the boldness indicator. As shown in columns (1)-(4), information from analysts who issue bold forecasts explains higher return volatility, supporting the argument of Clement and Tse (2005) that analysts provide bold forecasts when they possess relevant private information. In contrast, the coefficient estimates in columns (5)-(8), which examine the relationship between boldness and price impact, are statistically insignificant.
@@ 3.3.3 The Value of Analyst Information around Earnings Announcements
Previous evidence underlines the information content of analyst reports in the first week following earnings announcements. Accordingly, I evaluate the information value of analyst reports issued promptly following earnings announcements and beyond that.
@@
To formally validate the observed pattern, I conduct a regression analysis of log information value on an indicator variable $Week$ which takes the value of one if the report is released within one week following earnings announcements and its interaction with earnings announcement day trading volume ($TradingVolume$). This model is motivated by prior literature suggesting that analysts’ interpretative role becomes more prominent with the complexity of financial accounting information (Chen et al., 2010). The trading volume serves as a proxy for investor disagreement, following Banerjee (2011) and Beckmann et al. (2024). The interaction term ($Week\times TradingVolume$) captures how post-earnings timing and investor disagreement jointly affect analyst information value. Stock and year-fixed effects are included to control for time-invariant stock characteristics and market-wide trends.
@@
Panel B reveals significant increases in both log return variance and log price impact following earnings announcements, as evidenced by positive coefficients for the $Week$ variable. Notably, the increase in the information value of analyst reports following earnings announcements is primarily attributed to higher explained return volatility, rather than changes in price impact. These findings indicate that analyst reports provide more valuable insights in the first week after earnings announcements, especially during periods of high investor uncertainty. Investors value analysts’ ability to promptly interpret complex earnings information and make sense of the latest earnings numbers.
@@ A. Additional Figures
Figure A5: Shapley Values for Topic Importance across Industries